OpenStudy (anonymous):

Factor each of the following by finding common monomial factors :) xy-3x 2x^2+3x-12xy

OpenStudy (anonymous):

From first there is x in both the terms: SO take out x common from it..

OpenStudy (anonymous):

What is the answer ?

OpenStudy (anonymous):

You tell me..

OpenStudy (anonymous):

But i dont know yet :(

OpenStudy (anonymous):

See,, I give you other example: ab - 3a = a(b -3) see this and compare with the given question and try to take common..

OpenStudy (anonymous):

Ok ! Thanks :)

OpenStudy (anonymous):

Like this ? The answer of your given equation is : (b-3(b-3) Iam Right ?

OpenStudy (anonymous):

What is your first equation?

OpenStudy (anonymous):

I Take out only their Commons . So what is your answer ?

OpenStudy (anonymous):

Tell me your answer I will tell you mine.. Ladies first..

OpenStudy (anonymous):

(b-3(b-3) Iam Right ?

OpenStudy (anonymous):

I am saying what is your first equation?? Check the question and then tell me..

OpenStudy (anonymous):

I do not sure with my answer :(

OpenStudy (anonymous):

You are not listening what I am asking to you..

OpenStudy (unklerhaukus):

\[xy-3x=x\times y-3\times x=x(y-3)\]

OpenStudy (anonymous):

my answer is x(y-3)

OpenStudy (anonymous):

Yes it is right then what is the problem in doing that??

OpenStudy (anonymous):

Nothing :)

OpenStudy (unklerhaukus):

\[2x^2+3x-12xy\]\[=(2\times x\times x)+(3\times x)-(12\times x\times y)\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (unklerhaukus):

so can you now again factor out a factor of x from those brackets ?

OpenStudy (anonymous):

Yeah :)

OpenStudy (unklerhaukus):

that did you get/

OpenStudy (anonymous):

Thanks :) I get IT !

OpenStudy (zzr0ck3r):

please stop messaging me with your every problem ty.